OpenStudy (anonymous):

Find the inverse of the function and state whether the inverse is a function Y=x^3-4

OpenStudy (anonymous):

To do inverse of functions you first switch the y and the x. So waht would your formula look like?

OpenStudy (anonymous):

x=y^3-4?

OpenStudy (anonymous):

Correct. now you have to solve for y and that is your answer. Do you need help with that?

OpenStudy (anonymous):

yes please, i don't understand how I would do that

OpenStudy (anonymous):

Ok we need to isolate the y value to one side of the equation. Show me that please.

OpenStudy (anonymous):

x+4=y^3

OpenStudy (anonymous):

Good. Now we need to get y^3 reduced to y. Remember what ever we do to one side of the equation we must do to the other. Show me.

OpenStudy (anonymous):

do i square root each side? so \[\sqrt{x+4}\]=y?

OpenStudy (anonymous):

Not square root but cube root since (y^3). thus answer is \[y=\sqrt[3]{x+4}\]

OpenStudy (anonymous):

Good work. Remember inverse functions simply switch x and y then solve for y.